OpenStudy (anonymous):

HELP What is the simplified form of 1 over 5 x y times the fraction 10 x squared over 7 y ?

OpenStudy (anonymous):

Sorry, I really don't know how to solve this... Can you draw it out so i can see if I really can't solve it?

OpenStudy (anonymous):

ill put a pic

OpenStudy (anonymous):

Thanks. And I really need help on mine </3 Can someone help me?

OpenStudy (anonymous):

hmm. okay

OpenStudy (anonymous):

First, you need to multiply them together.

OpenStudy (mertsj):

\[\frac{1}{5xy}\times\frac{10x^2}{7y}\]

OpenStudy (anonymous):

^yup !

OpenStudy (mertsj):

Cancel the 5 and 10. Cancel the x. Multiply what is left.

OpenStudy (anonymous):

10x^2 ------ 35xy^2 2x ---- 7y^2

OpenStudy (mertsj):

\[\frac{2x}{7y^2}\]

OpenStudy (anonymous):

Yay! lol.